Lydia has money in two savings accounts. One rate is 7% and the other is 14%. If she has $650 more in the 14% account and the total interest is $133, how much is invested in each savings account?
Answer by mananth(16946)

Lydia has money in two savings accounts.
One rate is 7% . Let amount be x
and the other is 14%.
she has $650 more in the 14% account (x+650)

and the total interest is $133,

7%x + 14%(x+650) =133
multiply by 100
7x+14(x+650)=133*100
7x+14x+9100 =13300
21x = 4200
/21
x=200
x+650 = 850
$200 & $ 850 invested in each savings account